# CogniSNN: A First Exploration to Random Graph Architecture based Spiking Neural Networks with Enhanced Expandability and Neuroplasticity

### v1

@@ Abstract
Despite advances in spiking neural networks (SNNs) in numerous tasks, their architectures remain highly similar to traditional artificial neural networks (ANNs), restricting their ability to mimic natural connections between biological neurons. This paper develops a new modeling paradigm for SNN with random graph architecture (RGA), termed Cognition-aware SNN (CogniSNN). Furthermore, we improve the expandability and neuroplasticity of CogniSNN by introducing a modified spiking residual neural node (ResNode) to counteract network degradation in deeper graph pathways, as well as a critical path-based algorithm that enables CogniSNN to perform continual learning on new tasks leveraging the features of the data and the RGA learned in the old task. Experiments show that CogniSNN with re-designed ResNode performs outstandingly in neuromorphic datasets with fewer parameters, achieving $95.5\%$ precision in the DVS-Gesture dataset with only 5 timesteps. The critical path-based approach decreases $3\%$ to $5\%$ forgetting while maintaining expected performance in learning new tasks that are similar to or distinct from the old ones. This study showcases the potential of RGA-based SNN and paves a new path for biologically inspired networks based on graph theory.
@@ 1 Introduction
Spiking neural networks (SNNs), inspired by biological nervous systems, provide enhanced biological interpretability and energy efficiency over traditional artificial neural networks (ANNs). Recently, various prominent SNN models have been developed, such as Spike VGG  Rathi and Roy (2021), Spike ResNet  Fang et al. (2021), and Spike Transformer  Zhou et al. (2022). These models maintain the performance benefits of conventional ANNs while leveraging the intrinsic energy efficiency of SNNs, achieving improved accuracy on neuromorphic datasets.
Despite substantial advancements achieved by contemporary SNNs, their fundamental architecture continues to parallel the computational graph framework of ANNs  Xie et al. (2017, 2019). This framework diverges considerably from the brain’s biological configuration, which resembles a random graph architecture (RGA). As early as the 1940s,  Turing (1948) likened randomly connected unorganized machines to an infant’s cerebral cortex, establishing the link between computational models and neuroscience.  Rosenblatt (1958) later suggested that the setup of each organism’s nervous system is primarily random and distinct at birth. The research by  Varshney et al. (2011); Watts and Strogatz (1998) showed that the nematode nervous system is a random, small-world graph with about 300 neurons, which further supports the RGA of the biological system. Besides, it is crucial to recognize that if SNN still inherits the technological framework of conventional ANN, it may encounter similar issues related to hyperparameter configuration in ANNs, such as the number of layers, leading to performance barriers. These insights urge a re-evaluation of the SNN framework through the RGA perspective.
@@
In contrast, employing RGA in the architecture design of SNNs could be a potential solution to the aforementioned bottleneck by using graph theory to account for the spatial features of neural network topology. In practice, a similar effort was made by Xie et al. who performed experiments to validate that RGA-based ANNs achieved equivalent performance to computational graph-based ANN in multiple datasets  Xie et al. (2019). To our knowledge, the most related evidence comes from the investigation conducted by Yan et al.  Yan et al. (2024) who performed a neural architecture search on SNN models, indirectly highlighting the promise of RGA-based SNNs. Despite the lack of direct studies on the feasibility of RGA-based SNNs, such intuition can be derived by comparing the phase of spike processing over all neurons in an SNN to graph signal processing, which is highly dependent on the graph structure.
@@
Apart from simply evaluating the feasibility of RGA-based SNNs, it is eager to explore the potential of RGA-based SNNs in exhibiting expandability and neuroplasticity, which are two fundamental characteristics inherent in the architecture of the human brain, yet challenging to implement in deep learning models. Specifically, expandability refers to the ability to expand the graph structure to accommodate task learning. Traditional computational graph-based deep learning models also face expandability such that a model with deep layers might not perform as well as a model with shallow layers. However, this problem is even more challenging for RGA-based SNNs because of two aspects. First, a long path in the graph also leads to gradient vanishing/exploding. Although ResNet-inspired SNN approaches Zhou et al. (2024) could address this problem, they require significant real-valued computations, undermining the biologically inspired nature of SNNs. Moreover, unlike computational graphs where connections exist solely between neurons from adjacent layers, an edge in an RGA-based SNN possibly links neurons, in analogy with the computational graph, across disparate layers. The situation is even worse when the scale of the graph is large. Consequently, it results in feature dimension misalignment at a neuron, rendering pooling mechanisms ineffective.
Neuroplasticity  Demarin and Morović (2014) refers to the capacity of an RGA-based SNN to rapidly reuse or modify the architecture to adapt to new tasks and avoid catastrophic forgetting. Drawing on strategies from ANN models, methods such as lifelong learning are viable ways to cope with this problem in computational graph-based ANNs and SNNs. More precisely, strategies to achieve this goal mainly include methods based on regularization Li and Hoiem (2017), replay techniques Chaudhry et al. (2021), architectural adjustments Abati et al. (2020), and knowledge distillationGou et al. (2021). In particular, methodologies that emphasize path adjustment, such as PathNet Imai et al. (2020), exhibit greater alignment with CogniSNN due to the absence of layer structure in RGA. However, implementing these methods in new tasks often requires searching for optimal unused paths, thus neglecting previously established paths. Moreover, the management of these paths becomes challenging due to task interference, a difficulty exacerbated when there is a significant domain shift between the source (old) tasks and the target (new) tasks.
@@
This study introduces a cognition-aware SNN model (CogniSNN), conceived with the notion that the SNN itself possesses expandability and neuroplasticity cognition. We utilize Erdős–Rényi (ER) and Watts–Strogatz (WS) graph frameworks to structure CogniSNN, on which we develop a continual learning algorithm based on critical paths. The main contributions are as follows.
@@
  • Expandability: An OR Gate and a tailored pooling mechanism are proposed, where the former addresses network degradation and non-spike transmission, whereas the latter corrects feature dimension misalignment.
@@
  • Neuroplasticity: We develop a critical path-focused method to improve the continual learning ability of RGA-based SNNs. By utilizing betweenness centrality to identify essential pathways, CogniSNN selectively retrains the parameters associated with these pathways, mitigating catastrophic forgetting while maintaining performance on new tasks.
  • Extensive experiments demonstrate that CogniSNN improves accuracy by approximately $4.8\%$, $2.2\%$, and $2.6\%$ across three datasets. Furthermore, our critical path-based continual learning algorithm allows CogniSNN to manage similar and different tasks with effective suppression of catastrophic forgetting by approximately $3.6\%$ and $5.1\%$, respectively.
@@ 2.1 Residual Structures in SNNs
Residual learning He et al. (2016a) is necessary for training large-scale deep neural networks. Spiking ResNet Hu et al. (2021) introduces the residual learning in SNNs for the first time, converting trained ResNet to a SNN by replacing the ReLU layers with spike neurons. On the one hand, although it is better than SNNs without residual operation, the information of residual blocks is not effectively preserved due to the activation of spike neurons after the residual operation. Therefore, when the number of layers exceeds 18, deep Spiking ResNet begins to degrade gradually. On the other hand, it does not verify whether directly trained SNN can address the degradation problem. To solve these two problems, SEW-ResNet Fang et al. (2021) places the residual operation after the spike neuron activation, and for the first time, directly trains a 152-layer SNN. However, due to the addition operation, output of the residual block is changed from spike form to integer form, which violates the principle of spike computation in SNNs, and leads to greater energy consumption. Moreover, the output is amplified infinitely as the layers deepen, leading to an extremely unstable training and computation process. MS-ResNet Hu et al. (2024) places spike neuron activation in front of the convolutional layer to minimize non-spike computation in the convolutional layer. However, the output of the residual block is still non-spike, which leads to floating-point signal transmission in the subsequent residual branching.
@@ 2.2 Continual Learning in SNNs
Continual learning enables networks to continually learn like the human brain and adapt to dynamic environments in the real world, with the same brain-like perspective as SNNs. Recently, researchers have begun to focus on continual learning in SNNs. Antonov et al. (2022) propose a novel regularization method based on synaptic noise and Langevin dynamics, to preserve important parameters for the locally STDP-trained SNNs in continual learning. Proietti et al. (2023) achieve continual learning by using experience replay on SNNs for the first time, and perform well in class-incremental learning and task-free continual learning scenarios. Inspired by structure of the human brain and learning process in the biological systems, Han et al. (2023a) propose a self-organizing regulation network that selectively activates different sparse pathways according to different tasks to learn new knowledge. Similarly, Shen et al. (2024) utilize trace-based K-Winner-Take-All and variable threshold components to achieve selective activation of sparse neuron populations, not only in the spatial dimension but also in the temporal dimension for continual learning under specific tasks. Han et al. (2023b) dynamically change the structure of SNNs to improve learning ability of the SNNs for new tasks by randomly growing new neurons and adaptively pruning redundant neurons.
Restricted by structure, some of the above approaches are not biologically plausible enough, some require extra memory usage, and some necessitate introduction of the complex selection mechanisms at the SNN layer, which greatly increases the complexity of SNN implementation. To better align with the memory processes of the human brain, we choose Learning Without Forgetting (LWF) Li and Hoiem (2017) as the basic method.
@@ 3.1 Spiking Neuron
SNNs employ neurons characterized by biological dynamics to more closely emulate brain functionality. The spiking generation in such neurons is typically divided into three stages: charging, discharging, and resetting. Upon receiving an input current, a spiking neuron accumulates it as membrane potential. The charging mechanism is mathematically represented by (1).
@@
where $H[t]$ is the membrane potential before discharging at moment $t$, $C[t]$ refers to the input current at moment $t$, $V[t-1]$ is the membrane potential after discharging at moment $t-1$, and $f$ varies with the configuration of neurons.
At any given time $t$, the output of a discharging process at a neuron depends on whether the membrane potential exceeds a threshold $V_{thr}$. This process is expressed as (4):
$S[t]=\Theta\left(H[t]-V_{thr}\right)=\left\{\begin{array}[]{ll}1,&H[t]\geq V_{thr}\\ 0,&H[t]<V_{thr}\end{array}\right.,$ (4)
where $S[t]$ is the output at $t$, and $\Theta$ is the Heaviside step function.
Following the discharging process, the membrane potential experiences the resetting phase, which is categorized into soft and hard resetting. In this study, we employ the soft resetting approach, and this mechanism is represented by (5).
@@ 3.2 Learning Without Forgetting
LWF is a continual learning approach without using old task data, and aims to solve catastrophic forgetting of old task while learning new task.
Specifically, the method first computes the soft target $Y_{o}$ using new task data $X$, shared parameters $\theta_{s}$ and old task specific parameters $\theta_{o}$, where both parameters come from a well-trained model on old task. The process is shown in Eq. (6).
@@
During training of the new task, the LWF method requires using $\hat{\theta}_{o}$ and new task specific parameters $\hat{\theta}_{n}$ being trained to obtain old task output $Y{}^{\prime}_{o}$ and new task output $Y{}^{\prime}_{n}$, respectively. These two processes are shown in Eq. (7).
@@
Finally, a total loss function $Loss$ is utilized to balance learning and forgetting, defined as follows:
$Loss=\lambda L_{old}(Y_{o},Y{}^{\prime}_{o})+L_{new}(Y_{n},Y{}^{\prime}_{n})+R(\hat{\theta}_{s},\hat{\theta}_{o},\hat{\theta}_{n}).$ (8)
The $L_{new}$, computed using new task output $Y{}^{\prime}_{n}$ and new task label $Y_{n}$, is used to ensure that the model learns the new task. Additionally, $L_{old}$, computed using the soft target $Y_{o}$ and the old task output $Y{}^{\prime}_{o}$, is used to mitigate the model’s forgetting of old tasks. $R$ is the regularization term introduced to prevent overfitting.
@@ 4 Modeling and Methodology
Figure 1: The overall framework of the Cognition-aware Spiking Neural Network(CogniSNN) with 7 nodes.
This study first concentrates on the implementation of the expandability of RGA-based SNNs by introducing CogniSNN, facilitated by ResNode and a tailored pooling strategy. Subsequently, we detail CogniSNN’s capability to manage diverse tasks through a novel critical-path continual learning approach.
@@ 4.1 Modeling of CogniSNN
The triplet component comprises a convolution layer (Conv), a batch normalization layer (BN), and a spiking neuron layer (SN). In the following, we will denote this triplet by $ConvBNSN$.
The RGA-based module is modeled as a directed acyclic graph using WS or ER generators referenced in Xie et al. (2019). With $N$ denoted as the number of nodes, the graph is formulated as $G=\{\mathcal{V},E\}$ where $\mathcal{V}$ is the set of nodes with $|\mathcal{V}|=N$, and $E$ is the set of edges. As depicted in the blue block of Fig.1, each node $v_{i}\in\mathcal{V}$ represents a ResNode, a variant of residual block used in  Hu et al. (2021). The design of ResNode attempts to solve the issue that a long propagation easily leads to the gradient vanishing or explosion. The adjacency matrix of $G$ is $A$, in which $A_{i,j}=w,w\neq 0$ indicates that there exists an edge from $v_{i}$ to $v_{j}$ with weight $w$. Otherwise, $A_{i,j}=0$. At time $t$, the input $I_{j}[t]$ to node $v_{j}$ from its predecessors $\mathcal{P}_{j}$ is given by the following:
@@
where $O_{i}[t]$ is the output of $v_{i}$, $\sigma(\cdot)$ is the sigmoid function, $TP$ and $SP$ stands for the tailored pooling and standard pooling. To avoid feature loss, $SP$ is defined as follows:
$SP(O_{i}[t])=\left\{\begin{array}[]{ll}AvgP(O_{i}[t]),&D(O_{i}[t])\geq\eta\\ O_{i}[t],&Otherwise\end{array}\right.$ (12)
where $AvgP(\cdot)$ denotes average pooling, $D(\cdot)$ denotes the dimension of a feature map and $\eta$ is a manually configured threshold set by 32 according to our experimental results. $TP(\cdot)$ aims to solve the issue of feature misalignment caused by random connections of the spike neuron nodes in RGA. The solutions to ResNode modeling and the tailored pooling mechanism are the keys to improving the expandability of CogniCNN. Further elaboration is provided below.

### v2

# CogniSNN: An Exploration to Random Graph Architecture based Spiking Neural Networks with Enhanced Depth-Scalability and Path-Plasticity

@@ Abstract
Currently, most spiking neural networks (SNNs) still mimic the chain-like hierarchical architecture in traditional artificial neural networks (ANNs). This method significantly differs from random connections between neurons found in biological brains, limiting the ability to model the evolving mechanisms of neural pathways in biological neural systems, particularly in terms of dynamic depth-scalability and adaptive path-plasticity. This paper develops a new modeling paradigm for SNNs with random graph architecture (RGA), termed Cognition-aware SNN (CogniSNN). Furthermore, we model the depth-scalability and path-plasticity in CogniSNN by introducing a modified spiking residual neural node (ResNode) to counteract network degradation in deeper graph pathways, as well as a critical path-based algorithm that enables CogniSNN to perform path reusability on new tasks leveraging the features of the data and the RGA learned in old tasks. Experiments show that the performance of CogniSNN with redesigned ResNode is comparable, even superior, to current state-of-the-art SNNs on neuromorphic datasets. The critical path-based approach effectively achieves path reuse capability while maintaining expected performance in learning new tasks that are similar to or distinct from the old ones. This study showcases the potential of RGA-based SNNs and paves a new path for modeling the fusion of computational neuroscience and deep intelligent agents. The code is available at github.com/Yongsheng124/CogniSNN.
[…]
123
@@ 1 Introduction
Spiking neural networks (SNNs), inspired by biological nervous systems, provide enhanced biological interpretability and energy efficiency over traditional artificial neural networks (ANNs). Recently, various prominent SNN models have been developed, such as Spiking VGG [2022_feng_multi], Spiking ResNet [2021_fang_deep], and Spiking Transformer [2022_zhou_spikformer]. These models maintain the performance benefits of conventional ANNs while leveraging the intrinsic energy efficiency of SNNs, achieving improved performance on neuromorphic datasets. However, these developments remain predominantly driven by deep learning paradigms, which significantly deviate from the original brain-inspired purpose of SNNs.
The human brain constructs extremely deep neural pathway networks through random connections among hundreds of billions of neurons, and this topology endows the nervous system with strong adaptability to process complex tasks [2023_shen_esl, 2023_han_adaptive]. When encountering new task challenges, the brain always selectively and frequently activates specific pathways, enabling a seamless transition from simple task processing to complex reasoning [2021_barron_neural, 2023_han_adaptive]. In this paper, we define these two fundamental characteristics as depth-scalability (enhancing network capacity through pathway depth extension) and path-plasticity (enabling lifelong learning through dynamic path selection), which collectively constitute the foundation of biological intelligence’s adaptive capabilities. Empirical evidence demonstrates these two mechanisms enable the brain to accomplish environmental adaptation and lifelong learning with ultra-low power consumption. However, existing SNNs have yet to effectively model these properties, whose architectures remain constrained by accuracy-oriented ANN paradigms, neglecting biologically validated efficient learning mechanisms. This divergence results in SNNs’ deficiencies in biological interpretability, energy efficiency, and dynamic environmental adaptability. We believe that developing a novel computational framework integrating depth-scalability with path-plasticity represents a critical breakthrough to advance brain-like intelligence.
@@
Current SNNs mainly adopt conventional chain-like hierarchical architectures [2017_xie_aggregated], making it difficult to effectively model neural pathways. In contrast, the brain’s biological structure resembles a random graph architecture (RGA) [1958_rosenblatt_perceptron, 2011_varshney_structural], and the modeling of neural pathways also requires the assistance of graph structures. Besides, it is crucial to recognize that if SNNs still inherit the technological framework of conventional ANN, it may encounter similar issues related to hyperparameter configuration in ANNs, such as the number of layers, leading to performance barriers [2023_xu_unified]. These insights urge a re-evaluation of the SNN framework through the RGA perspective. In practice, a similar effort was made by Xie et al., who performed experiments to validate that RGA-based ANNs achieved equivalent performance to computational graph-based ANNs in multiple datasets [2019_xie_exploring]. To our knowledge, the most relevant evidence comes from the investigation conducted by Yan et al., who performed a neural architecture search on SNN models [2024_yan_sampling], indirectly highlighting the promise of RGA-based SNNs. Some studies [2021_xu_exploiting, 2023_xu_exploiting] combine graph learning and SNNs, but mainly focus on graph data rather than graph structure of the network. Despite the lack of direct studies on the feasibility of RGA-based SNNs, such intuition can be derived by comparing the phase of spike processing over all neurons in an SNN to graph signal processing, which is highly dependent on the graph structure.
@@
After simply evaluating the feasibility of RGA-based SNNs, it is eager to explore the potential of RGA-based SNNs in exhibiting depth-scalability and path-plasticity, which are two fundamental characteristics inherent in the architecture of the human brain, yet challenging to implement in RGA-based SNN models. Specifically, in terms of depth-scalability, first, a long path in the graph leads to gradient vanishing/exploding. Although ResNet-inspired SNN approaches could address this problem, they require significant real-valued computations [2024_zhou_direct], undermining the biologically inspired nature of SNNs. Moreover, unlike computational graphs where connections exist solely between neurons from adjacent layers, an edge in an RGA-based SNN possibly links neurons across disparate layers. Consequently, it results in feature dimension misalignment, rendering pooling mechanisms ineffective.
In terms of path-plasticity, drawing from continual learning technologies, strategies to achieve this goal mainly include methods based on regularization [2024_xu_adaptive], replay techniques [2021_chaudhry_using], architectural adjustments [2025_shen_context], and knowledge distillation [2025_li_frequency]. However, due to the limitations of conventional neural network architectures, similar architectural adjustment methods tend to focus on the internal structure of neurons within convolutional layers, neglecting the path structures in graph-structured networks. This makes these methods unsuitable for RGA-based SNNs. Moreover, the management of these paths becomes challenging due to task interference, a difficulty exacerbated when there is a significant domain shift between the source (old) tasks and the target (new) tasks.
@@
This study introduces a Cognition-aware SNN model (CogniSNN), conceived with the notion of efficiently modeling the depth-scalability and path-plasticity cognition that is inherently present in the biological brain through RGA-based SNNs. We utilize Erdős–Rényi (ER) and Watts–Strogatz (WS) graph frameworks to structure CogniSNN, on which we develop a continual learning algorithm based on critical paths. The main contributions are as follows.
@@
  • Depth-Scalability: An OR Gate and a tailored pooling mechanism are proposed, where the former addresses network degradation and non-spike transmission, whereas the latter corrects feature dimension misalignment.
@@
  • Path-Plasticity: We develop a critical path-focused method to improve the pathway reuse capability of RGA-based SNNs. By utilizing betweenness centrality to identify essential pathways, CogniSNN selectively retrains the parameters associated with these pathways, mitigating catastrophic forgetting while maintaining performance on new tasks.
  • Extensive experiments show that CogniSNN achieves performance comparable to or even surpassing current state-of-the-art models on all three datasets. In addition, our critical path-based path adaptive mechanism enables CogniSNN to manage both similar and dissimilar tasks, effectively suppressing catastrophic forgetting.
@@ 2.1 Residual Structures in SNNs
Residual learning [2016_he_deep] is necessary for training large-scale deep neural networks. Spiking ResNet [2021_hu_spiking] introduces residual learning in SNNs for the first time, converting trained ResNet to an SNN by replacing ReLU layers with spike neurons. On the one hand, although it is better than SNNs without residual operation, the information of residual blocks is not effectively preserved due to the activation of spike neurons after the residual operation. Therefore, when the number of layers exceeds 18, deep Spiking ResNet begins to degrade gradually. On the other hand, it does not verify whether directly trained SNNs can address the degradation problem. To solve these two problems, SEW-ResNet [2021_fang_deep] places the residual operation after the spike neuron activation, and for the first time, directly trains a 152-layer SNN. However, due to the addition operation, output of the residual block is changed from spike form to integer form, which violates the principle of spike computation in SNNs, and leads to greater energy consumption. Moreover, the output is amplified infinitely as the layers deepen, leading to an extremely unstable training and computation process. MS-ResNet [2024_hu_advancing] places spike neuron activation in front of the convolutional layer to minimize non-spike computation in the convolutional layer. However, the output of the residual block is still non-spike, which leads to floating-point signal transmission in the subsequent residual branching.
@@ 2.2 Continual Learning in SNNs
Continual learning enables networks to continually learn as the human brain and adapt to dynamic environments in the real world, with the same brain-like perspective as SNNs. Recently, researchers have begun to focus on continual learning in SNNs. [2022_antonov_continuous] propose a novel regularization method based on synaptic noise and Langevin dynamics, to preserve important parameters for the locally STDP-trained SNNs in continual learning. [2023_proietti_memory] achieve continual learning by using experience replay on SNNs for the first time, and perform well in class-incremental learning and task-free continual learning scenarios. Inspired by the structure of the human brain and learning process in the biological system, [2023_han_adaptive] propose a self-organizing regulation network that selectively activates different sparse pathways according to different tasks to learn new knowledge. Similarly, [2024_shen_efficient] utilize trace-based K-Winner-Take-All and variable threshold components to achieve selective activation of sparse neuron populations, not only in the spatial dimension but also in the temporal dimension for continual learning under specific tasks. [2023_han_enhancing] dynamically change the structure of SNNs to improve learning ability of the SNNs for new tasks by randomly growing new neurons and adaptively pruning redundant neurons.
Restricted by structure, some of the above approaches are not biologically plausible enough, some require extra memory usage, and some necessitate the introduction of the complex selection mechanisms at the spiking neuron layer, which greatly increases the complexity of SNN implementation.
@@ 3.1 Spiking Neuron
SNNs employ neurons characterized by biological dynamics to more closely emulate brain functionality. The generation of spike in such neurons is typically divided into three stages: charging, discharging, and resetting. Upon receiving an input current, a spiking neuron accumulates it as membrane potential. The charging mechanism is mathematically represented by:
@@
where $H[t]$ is the membrane potential before discharging at moment $t$, $C[t]$ refers to the input current at moment $t$, V[t1]V[t-1] is the membrane potential after discharging at moment t1t-1, and $f(\cdot)$ varies with the configuration of neurons.
At any given time $t$, the output of a discharging process at a neuron depends on whether the membrane potential exceeds a threshold $V_{thr}$. This process is expressed as:
$S[t]=\Phi\left(H[t]-V_{thr}\right)=\left\{\begin{array}[]{ll}1,&H[t]\geq V_{thr}\\ 0,&H[t]<V_{thr}\end{array}\right$., (4)
where $S[t]$ is the output at $t$, and $\Phi$ is the Heaviside step function.
Following the discharging process, the membrane potential experiences the resetting phase, which is categorized into soft and hard resetting. In this study, we employ the soft resetting approach, and this mechanism is represented by:
@@ 3.2 Learning without Forgetting
Learning without forgetting (LwF) [2017_li_learning] is a continual learning approach without using old task data, and aims to solve catastrophic forgetting of old task while learning new task.
Specifically, the method first computes the soft target $Y_{o}$ using new task data $X$, shared parameters $\theta_{s}$ and old task specific parameters $\theta_{o}$, where both parameters come from a well-trained model on old task. The process is mathematically represented by:
@@
During training of the new task, the LwF method requires using $\hat{\theta}_{o}$ and new task specific parameters $\hat{\theta}_{n}$ being trained to obtain old task output $Y^{\prime}_{o}$ and new task output $Y^{\prime}_{n}$, respectively. These two processes are shown by:
@@
Finally, a total loss $\mathcal{L}$ is utilized to balance learning and forgetting, defined as follows:
$\mathcal{L}=\lambda L_{old}(Y_{o},Y^{\prime}_{o})+L_{new}(Y_{n},Y^{\prime}_{n})+R(\hat{\theta}_{s},\hat{\theta}_{o},\hat{\theta}_{n}).$ (8)
The $L_{new}(\cdot)$, computed using new task output $Y^{\prime}_{n}$ and new task label $Y_{n}$, is used to ensure that the model learns the new task. Additionally, $L_{old}(\cdot)$, computed using the soft target $Y_{o}$ and the old task output $Y^{\prime}_{o}$, is used to mitigate the model’s forgetting of old tasks. $R(\cdot)$ is the regularization term introduced to prevent overfitting.
@@ 4 Modeling and Methodology
Figure 1: Comprehensive diagram of CogniSNN. During neuromorphic object recognition, data flows from the ConvBNSNConvBNSN triple to the RGA-based SNN and then enters the classifier. The internal structure of ResNode and the illustration of dimension misalignment are shown at the bottom right and right side of the figure, respectively. During continual learning, different critical paths are selected for learning new knowledge in different scenarios.
This study first concentrates on the implementation of the depth-scalability of RGA-based SNNs by introducing CogniSNN, facilitated by ResNode and a tailored pooling strategy. Subsequently, we detail CogniSNN’s capability to manage diverse tasks through a novel critical-path continual learning approach.
@@ 4.1 Modeling of CogniSNN
The triplet component comprises a convolution layer (Conv), a batch normalization layer (BN), and a spiking neuron layer (SN). In the following, we denote this triplet by $ConvBNSN$.
The RGA-based module is modeled as a directed acyclic graph using WS or ER generators referenced in [2019_xie_exploring]. With $N$ denoted as the number of nodes, the graph is formulated as $G=\{\mathcal{V},E\}$, where $\mathcal{V}$ is the set of nodes with $|\mathcal{V}|=N$, and $E$ is the set of edges. As depicted in the yellow block of Fig.1, each node $v_{i}\in\mathcal{V}$ represents a ResNode, a variant of the residual block used in [2021_hu_spiking]. The design of ResNode attempts to solve the issue that a long propagation easily leads to the gradient vanishing or explosion. The adjacency matrix of $G$ is $A$, in which $A_{i,j}=w,w\neq 0$ indicates that there exists an edge from $v_{i}$ to $v_{j}$ with weight $w$. Otherwise, $A_{i,j}=0$. At time $t$, the input $I_{j}[t]$ to node $v_{j}$ from its predecessors $\mathcal{P}_{j}$ is given by the following:
@@
where $O_{i}[t]$ is the output of $v_{i}$, $\sigma(\cdot)$ is the sigmoid function, $TP$ and $SP$ stand for the tailored pooling and standard pooling. To avoid feature loss, $SP$ is defined as follows:
$SP(O_{i}[t])=\left\{\begin{array}[]{ll}AvgP(O_{i}[t],\kappa),&D(O_{i}[t])\geq\eta\\ O_{i}[t]&Otherwise\end{array},\right$. (12)
where $AvgP(\cdot)$ denotes average pooling, $D(\cdot)$ denotes the spatial dimension of a feature map, and $\eta$ is set to 1 and 32 in the classification and continual learning experiments, respectively. $\kappa$ represents the pooling kernel size and is generally set to 2. $TP(\cdot)$ aims to solve the issue of feature misalignment caused by random connections of the spike neuron nodes in RGA. The solutions to ResNode modeling and the tailored pooling mechanism are the keys to improving the depth-scalability of CogniSNN. Further elaboration is provided below.